how do you solve 1/3p=2/5p+2/3
Answer by stanbon(48559)

1/3p=2/5p+2/3
------------
Multiply thru by 15p to get:
5 = 6 + 10p
10p = -1
p = -1/10
=============
Cheers,
Stan H.